OpenStudy (anonymous):

The ratio of the numerator to the denominator of a certain fraction is three to five. If two is added to the numerator and five is subtracted from the denominator, the new fraction reduces to four-fifths. Find the original fraction. 3/5 20/25 18/30

OpenStudy (whpalmer4):

Initial fraction is reducible to \[\frac{3}{5}\] That means we can write it as \[\frac{3x}{5x}\]where \(x\) is some unknown factor. Do you agree that expression gives us the correct ratio?

OpenStudy (anonymous):

Let x and y be the numerator and the denominator of the initial fraction, respectively. Solve the following equations for x and y.\[\left\{\frac{x}{y}=\frac{3}{5},\frac{x+2}{y-5}=\frac{4}{5}\right\} \]x=18 and y=30 \[\frac{18}{30}=\frac{3}{5} \]and\[\frac{18+2}{30-5}=\frac{4}{5} \]

OpenStudy (whpalmer4):

My approach doesn't require solving a system of equations, but is otherwise equivalent. \[\frac{3x+2}{5x-5} = \frac{4}{5}\]Cross-multiply and solve for \(x\): \[5(3x+2) = 4(5x-5)\]\[15x+10=20x-20\]\[30=5x\]\[x=6\]therefore, fraction is \[\frac{3(6)}{5(6)} = \frac{18}{30}\]